OpenStudy (anonymous):

5s=7s + 1 - 2s The book says the anwser is 9?

OpenStudy (anonymous):

what is the value of s?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

value = what number does s mean? 1,3, 7 , 10 ?? 5s means 5 x s.

OpenStudy (anonymous):

when i do the math i get 1

OpenStudy (anonymous):

so 7 x 1 = 7 + 1 = 8 - 2 x1 = 4?? But 5x 1 = 5. I think the number is off by 1.

OpenStudy (anonymous):

I tryied 9 is what the book says but i get 1 number off on it.

OpenStudy (anonymous):

so this is unsolveable?

OpenStudy (anonymous):

I think it might be a textbook error. It happens sometimes. If you cannot solve accurately, make sure to inform your teacher.

OpenStudy (anonymous):

I will thank you

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

only been working the problem for a hour and getting frusted